OpenStudy (anonymous):

0.8 = 8/10 True Fals

OpenStudy (anonymous):

True

OpenStudy (anonymous):

-0.28764 . . . is: rational irrational

OpenStudy (unklerhaukus):

\[0.8=0.8 \times1=0.8\times\frac{10}{10}=\frac{0.8\times10}{10}=\frac{8}{10}\]

OpenStudy (anonymous):

Does the "..." after the the number mean that it just keeps going? Because if so, it's an irrational number.

OpenStudy (unklerhaukus):

-0.287644444444444444444... is rational -0.287642876428764287642... is rational -0.287647728390198872933... looks like its going to be irrational because the it dosent look like its forming a pattern